# Long-term impact of PM2.5 on mortality is exacerbated when wildfire events occur

### v1

@@ 1 Main
There is extensive and consistent evidence that long-term exposure to PM2.5 from all sources has an adverse effect on human health and increases the risk of mortality [1, 2, 3, 4, 5, 6]. This robust evidence base led to a recent revision of the annual standard for the US National Ambient Air Quality Standards (NAAQS) for PM2.5, reducing the allowable threshold from 12 $\mu g/m^{3}$ to 9 $\mu g/m^{3}$. These more stringent standards are expected to yield significant public health benefits, including the prevention of up to 4,500 premature deaths per year across the US. Moreover, the net economic benefits—accounting for reduced healthcare expenditures, fewer workdays lost, and other health-related gains minus the estimated costs of regulatory compliance—are projected to reach approximately \$46 billion by 2032 [7].
The increasing frequency and geographic scope of wildfires and their smoke have increased interest in their health impacts. Climate-induced increases in fuel aridity, historical fire suppression, and the growing human presence in wildland areas have driven increased wildfire activity.[8] Since 2000, an average of 70,025 wildfires burned 7.0 million acres annually in the US, more than twice the average of the 1990s [9, 10]. Beginning in 2016, these events have reversed decades of progress in reducing PM2.5 concentrations in nearly three-quarters of US states [11].
Several studies have reported an association between short- and medium-term exposure to wildfire events and higher rates of emergency department visits, hospitalizations, and consultations for respiratory and cardiovascular problems [12, 13, 14, 15], as well as a higher risk of all-cause mortality [16, 17, 18] and and deaths due to COVID-19 [19]. In these studies, wildfire events were defined either based on the levels of wildfire PM2.5–fine particulate matter originating from wildfire smoke–or all-source PM2.5 during periods when wildfires burned. Recent studies have documented associations between long-term exposure to wildfire PM2.5 and adverse respiratory, mental, and overall health outcomes, including an increased risk of mortality [20, 21, 22].
@@
Our study cohort comprises more than 60 million US adults aged 65 years and older, with a total of 17,608,624 recorded deaths from 2007 to 2016. Exposure to PM2.5 is characterized as the annual average concentration of all-source PM2.5 between 2006 and 2015, as utilized by Wu et al. [2], derived from the daily estimates developed by Di et al. [23]. The potential effect modification by wildfire events is defined as the ”number of days with non-zero wildfire PM2.5 per year” [24]. Estimates of daily wildfire PM2.5 are obtained from a previously validated daily model for the contiguous US over the period 2006–2015 [25]. This metric effectively captures both the duration and frequency of exposure to wildfire events in each year. For our analyses, we categorize this effect modifier into three distinct strata based on annual counts of wildfire event days: (1) 0 to 20 non-zero wildfire PM2.5 days per year (approximating the first tertile); (2) 21 to 35 non-zero wildfire PM2.5 days per year (approximating the second tertile); and (3) more than 35 non-zero wildfire PM2.5 days per year. Figure 1 illustrates the spatial and temporal distribution of these strata across the US. Exposure and effect modification variables are aggregated at the ZIP code level and linked to Medicare beneficiaries according to their residential ZIP code and year of assessment. This linkage enables us to associate mortality data with exposure measures and effect modifiers from the previous year.
Within each of the three wildfire event days strata, we estimate the association between annual all-source PM2.5 levels and all-cause mortality counts in the subsequent year for each ZIP code using Poisson regression models allowing for a nonlinear exposure–response relationship between all-source PM2.5 and mortality. These models are also stratified by individual-level characteristics and follow-up year. To control for additional potential confounding, we adjust for 10 ZIP code or county-level covariates, four ZIP code–level meteorological variables, and indicators for geographic region (Northeast, South, Midwest, West) and calendar year [2]. To examine additional potential effect modification by geography or socioeconomic status, we further stratify the cohort by region and ZIP code-level poverty, using 2010 US Census data. ZIP codes are categorized as high or low poverty based on whether 15% or more of the population lives below the federal poverty level. Geographic and poverty-related modifiers, along with potential confounders, were assigned for the year corresponding to the outcome.
Our main parameter of interest is the relative change in all-cause mortality risk associated with deviations from the current National Ambient Air Quality Standard (NAAQS) for annual PM2.5 levels (9 $\mu g/m^{3}$), under each strata (0–20, 21–35, and $>$35 wildfire PM2.5 days per year).
@@
In this study, we find that experiencing more days of non-zero wildfire PM2.5 amplifies the association between long-term exposure to all-source PM2.5 and all-cause mortality in a cohort of $>$60 million older US adults, providing a new and realistic conceptualization of wildfire health risks.
@@ A.2 Exposure assessment
For the period 2006-2015, we consider the annual average of all-source PM2.5 concentrations at the ZIP code level as the exposure derived from daily estimates by Di et al. [23]. These daily estimates were obtained from an ensemble prediction model [23]. We then obtained ZIP code level estimates by averaging grid cells with centroids inside each ZIP code. For P.O. Box-only ZIP codes, all-source PM2.5 concentrations were assigned based on the nearest grid cell. Finally, annual concentrations were obtained by averaging the daily levels. We assign exposure based on the Medicare enrollee’s residential ZIP code in each calendar year, linking the prior year’s exposure to the current year’s mortality outcome.
@@ A.3 Potential effect modifiers
For the period 2006-2015, we consider the number of days with non-zero wildfire PM2.5 per year as a potential effect modifier of the all-source PM2.5-mortality association. We obtain daily levels of wildfire PM2.5 at 10-km2 grid resolution from Childs et al. [25]. We overlay the 10-km2 grid and the ZIP code boundaries and use area-weighting to obtain daily ZIP code level wildfire PM2.5 estimates. For P.O. Box-only ZIP codes, wildfire PM2.5 concentrations are assigned based on the nearest non-P.O. Box ZIP code. We then use the daily ZIP code estimates of wildfire PM2.5 to compute the annual number of non-zero wildfire PM2.5 days for each ZIP code. Following [24], we count the number of days where wildfire PM2.5 concentrations exceed zero per year in each ZIP code. Lastly, we categorize the annual number of non-zero wildfire PM2.5 days in three strata: (1) 0 to 20 non-zero wildfire PM2.5 days per year (approximating the first tertile); (2) 21 to 35 non-zero wildfire PM2.5 days per year (approximating the second tertile); and (3) more than 35 non-zero wildfire PM2.5 days per year. We assign wildfire event days based on the Medicare enrollee’s residential ZIP code in each calendar year, linking the prior year’s number of days with non-zero wildfire PM2.5 to the current year’s mortality.
@@ A.4 Potential confounders
In order to address confounding bias from community-level factors, we use information from various sources, including ZIP code-level socioeconomic status (SES) indicators from the 2000 and 2010 Census and the 2005–2012 American Community Surveys (ACS) and county-level information from the Centers for Disease Control and Prevention’s Behavioral Risk Factor Surveillance System (BRFSS). We spatially align all data with postal ZIP codes to match Medicare resolution. This includes: (i) two county-level variables: average body mass index and smoking rate; (ii) eight ZIP census variables: proportion of Hispanic and black residents, median household income, median home value, proportion of residents in poverty, proportion of residents with a high school diploma, population density, and proportion of residents that own their house; and (iii) four ZIP code-level meteorological variables from Gridmet [26, 27]: summer (June to September) and winter (December to February) average maximum daily temperature and relative humidity. The ZIP code–level meteorological variables are obtained using area-weighted aggregations based on daily temperature and humidity data on 4-km2 gridded rasters from Gridmet via Google Earth Engine. Finally, we create categorical variables for the four geographic US census regions (Northeast, South, Midwest, and West), and calendar year (2007–2016), to help account for any residual or unmeasured spatial and temporal confounding. This set of confounders corresponds to the ones used in [2].
@@ A.5 Data linkage
Mortality data are available at the ZIP code level. ZIP codes that have incomplete data or no Medicare patients are excluded from the analysis (n = 9,882). The analysis includes 34,444 ZIP codes ($\sim$80of US ZIP codes) with complete data on outcomes, exposures, effect modifiers, and confounders.
@@ A.6 Statistical analysis
To assess this potential effect modification, we implement a stratified Poisson regression approach, where the study cohort is divided into three strata based on the number of days with $>$0 $\mu g/m^{3}$ of wildfire PM2.5 per year: (1) 0–20 days, (2) 21–35 days, and (3) more than 35 days. Within each wildfire day stratum, we independently estimate the exposure–response relationship between annual average all-source PM2.5 levels and mortality. This stratified design allows us to directly compare the hazard ratios for PM2.5 across different levels of wildfire days.
The Poisson model is specified for each stratum using ZIP code–level death counts by calendar year and follow-up year as the outcome, with annual average all-source PM2.5 as a time-varying exposure and the log of person-years included as an offset term. We incorporate natural spline terms to model non-linear associations with PM2.5. Models are adjusted for 10 ZIP code or county-level socioeconomic and demographic covariates, four ZIP code–level meteorological variables, as well as fixed effects for geographic region (Northeast, South, Midwest, West) and calendar year to account for geographic and temporal heterogeneity. To control for individual-level variability in baseline mortality risk, each model is stratified by four individual-level characteristics: (i) a 5-year category of age at entry (65 to 69, 70 to 74, 75 to 79, 80 to 84, 85 to 89, 90 to 94, 95 to 99, and above 100 years of age), (ii) race/ethnicity (Hispanic, Native American and Alaska Native, Non-Hispanic Asian, Non-Hispanic Black, Non-Hispanic White, and other [e.g., multi-racial and missing]), (iii) sex (male or female), and (iv) an indicator variable for Medicaid eligibility (a surrogate for individual-level SES). For more details, see [2].
@@
[…]
@@ Acknowledgments
Table S1: Characteristics for the Medicare study cohort, 2006–2016. Mortality and individual-level characteristics were obtained from the Centers for Medicare and Medicaid Services (CMS), and ZIP code socioeconomic status (SES) was obtained from the 2000 and 2010 Census and the 2005–2012 American Community Surveys (ACS) and county-level behavioral risk factor variables were obtained from the Centers for Disease Control and Prevention. All-source PM2.5 and wildfire PM2.5 concentrations were obtained respectively from [3] and [25]. Meteorological variables were obtained from Gridmet via Google Earth Engine.
@@
$>$15% $\leq$15Northeast Midwest South West

### v2

@@ 1 Main
There is extensive and consistent evidence that long-term exposure to PM2.5 from all sources has an adverse effect on human health and increases the risk of mortality [josey2023air, wu2020evaluating, di2017air, moon2024long, brunekreef2021mortality, pun2017long]. This robust evidence base led to a recent revision of the annual standard for the US National Ambient Air Quality Standards (NAAQS) for PM2.5, reducing the allowable threshold from 12 $\mu g/m^{3}$ to 9 $\mu g/m^{3}$. These more stringent standards are expected to yield significant public health benefits, including the prevention of up to 4,500 premature deaths per year across the US. Moreover, the net economic benefits—accounting for reduced healthcare expenditures, fewer workdays lost, and other health-related gains minus the estimated costs of regulatory compliance—are projected to reach approximately \$46 billion by 2032 [EPAfinal].
The increasing frequency and geographic scope of wildfires and their smoke have increased interest in their health impacts. Climate-induced increases in fuel aridity, historical fire suppression, and the growing human presence in wildland areas have driven increased wildfire activity.[parks2020warmer] Since 2000, an average of 70,025 wildfires burned 7.0 million acres annually in the US, more than twice the average of the 1990s [epa_wildfires, nifc_statistics]. Beginning in 2016, these events have reversed decades of progress in reducing PM2.5 concentrations in nearly three-quarters of US states [Burke2023].
Several studies have reported an association between short- and medium-term exposure to wildfire events and higher rates of emergency department visits, hospitalizations, and consultations for respiratory and cardiovascular problems [Chen2021, Groot2019, To2021, Reid2024], as well as a higher risk of all-cause mortality [Chen2021mortality, Ye2022, Doubleday2020] and and deaths due to COVID-19 [Zhou2021]. In these studies, wildfire events were defined either based on the levels of wildfire PM2.5–fine particulate matter originating from wildfire smoke–or all-source PM2.5 during periods when wildfires burned. Recent studies have documented associations between long-term exposure to wildfire PM2.5 and adverse respiratory, mental, and overall health outcomes, including an increased risk of mortality [Connolly2024, Ma2024, Wei2023].
@@
Our study cohort comprises more than 60 million US adults aged 65 years and older, with a total of 17,608,624 recorded deaths from 2007 to 2016. Exposure to PM2.5 is characterized as the annual average concentration of all-source PM2.5 between 2006 and 2015, as utilized by Wu et al. [wu2020evaluating], derived from the daily estimates developed by Di et al. [DI2019104909]. The potential effect modification by wildfire events is defined as the ”number of days with non-zero wildfire PM2.5 per year” [Casey2024]. Estimates of daily wildfire PM2.5 are obtained from a previously validated daily model for the contiguous US over the period 2006–2015 [childs2022daily]. This metric effectively captures both the duration and frequency of exposure to wildfire events in each year. For our analyses, we categorize this effect modifier into three distinct strata based on annual counts of wildfire event days: (1) 0 to 20 non-zero wildfire PM2.5 days per year (approximating the first tertile); (2) 21 to 35 non-zero wildfire PM2.5 days per year (approximating the second tertile); and (3) more than 35 non-zero wildfire PM2.5 days per year. Figure 1 illustrates the spatial and temporal distribution of these strata across the US. Exposure and effect modification variables are aggregated at the ZIP code level and linked to Medicare beneficiaries according to their residential ZIP code and year of assessment. This linkage enables us to associate mortality data with exposure measures and effect modifiers from the previous year.
Within each of the three wildfire event days strata, we estimate the association between annual all-source PM2.5 levels and all-cause mortality counts in the subsequent year for each ZIP code using Poisson regression models allowing for a nonlinear exposure–response relationship between all-source PM2.5 and mortality. These models are also stratified by individual-level characteristics and follow-up year. To control for additional potential confounding, we adjust for 10 ZIP code or county-level covariates, four ZIP code–level meteorological variables, and indicators for geographic region (Northeast, South, Midwest, West) and calendar year [wu2020evaluating]. To examine additional potential effect modification by geography or socioeconomic status, we further stratify the cohort by region and ZIP code-level poverty, using 2010 US Census data. ZIP codes are categorized as high or low poverty based on whether 15% or more of the population lives below the federal poverty level. Geographic and poverty-related modifiers, along with potential confounders, were assigned for the year corresponding to the outcome.
Our main parameter of interest is the relative change in all-cause mortality risk associated with deviations from the current National Ambient Air Quality Standard (NAAQS) for annual PM2.5 levels (9 $\mu g/m^{3}$), under each strata (0–20, 21–35, and >>35 wildfire PM2.5 days per year).
@@
In this study, we find that experiencing more days of non-zero wildfire PM2.5 amplifies the association between long-term exposure to all-source PM2.5 and all-cause mortality in a cohort of >>60 million older US adults, providing a new and realistic conceptualization of wildfire health risks.
@@ A.2 Exposure assessment
For the period 2006-2015, we consider the annual average of all-source PM2.5 concentrations at the ZIP code level as the exposure derived from daily estimates by Di et al. [DI2019104909]. These daily estimates were obtained from an ensemble prediction model [DI2019104909]. We then obtained ZIP code level estimates by averaging grid cells with centroids inside each ZIP code. For P.O. Box-only ZIP codes, all-source PM2.5 concentrations were assigned based on the nearest grid cell. Finally, annual concentrations were obtained by averaging the daily levels. We assign exposure based on the Medicare enrollee’s residential ZIP code in each calendar year, linking the prior year’s exposure to the current year’s mortality outcome.
@@ A.3 Potential effect modifiers
For the period 2006-2015, we consider the number of days with non-zero wildfire PM2.5 per year as a potential effect modifier of the all-source PM2.5-mortality association. We obtain daily levels of wildfire PM2.5 at 10-km2 grid resolution from Childs et al. [childs2022daily]. We overlay the 10-km2 grid and the ZIP code boundaries and use area-weighting to obtain daily ZIP code level wildfire PM2.5 estimates. For P.O. Box-only ZIP codes, wildfire PM2.5 concentrations are assigned based on the nearest non-P.O. Box ZIP code. We then use the daily ZIP code estimates of wildfire PM2.5 to compute the annual number of non-zero wildfire PM2.5 days for each ZIP code. Following [Casey2024], we count the number of days where wildfire PM2.5 concentrations exceed zero per year in each ZIP code. Lastly, we categorize the annual number of non-zero wildfire PM2.5 days in three strata: (1) 0 to 20 non-zero wildfire PM2.5 days per year (approximating the first tertile); (2) 21 to 35 non-zero wildfire PM2.5 days per year (approximating the second tertile); and (3) more than 35 non-zero wildfire PM2.5 days per year. We assign wildfire event days based on the Medicare enrollee’s residential ZIP code in each calendar year, linking the prior year’s number of days with non-zero wildfire PM2.5 to the current year’s mortality.
@@ A.4 Potential confounders
In order to address confounding bias from community-level factors, we use information from various sources, including ZIP code-level socioeconomic status (SES) indicators from the 2000 and 2010 Census and the 2005–2012 American Community Surveys (ACS) and county-level information from the Centers for Disease Control and Prevention’s Behavioral Risk Factor Surveillance System (BRFSS). We spatially align all data with postal ZIP codes to match Medicare resolution. This includes: (i) two county-level variables: average body mass index and smoking rate; (ii) eight ZIP census variables: proportion of Hispanic and black residents, median household income, median home value, proportion of residents in poverty, proportion of residents with a high school diploma, population density, and proportion of residents that own their house; and (iii) four ZIP code-level meteorological variables from Gridmet [abatzoglou2013development, gorelick2017google]: summer (June to September) and winter (December to February) average maximum daily temperature and relative humidity. The ZIP code–level meteorological variables are obtained using area-weighted aggregations based on daily temperature and humidity data on 4-km2 gridded rasters from Gridmet via Google Earth Engine. Finally, we create categorical variables for the four geographic US census regions (Northeast, South, Midwest, and West), and calendar year (2007–2016), to help account for any residual or unmeasured spatial and temporal confounding. This set of confounders corresponds to the ones used in [wu2020evaluating].
@@ A.5 Data linkage
Mortality data are available at the ZIP code level. ZIP codes that have incomplete data or no Medicare patients are excluded from the analysis (n = 9,882). The analysis includes 34,444 ZIP codes ($\sim$80% of US ZIP codes) with complete data on outcomes, exposures, effect modifiers, and confounders.
@@ A.6 Statistical analysis
To assess this potential effect modification, we implement a stratified Poisson regression approach, where the study cohort is divided into three strata based on the number of days with >>0 $\mu g/m^{3}$ of wildfire PM2.5 per year: (1) 0–20 days, (2) 21–35 days, and (3) more than 35 days. Within each wildfire day stratum, we independently estimate the exposure–response relationship between annual average all-source PM2.5 levels and mortality. This stratified design allows us to directly compare the hazard ratios for PM2.5 across different levels of wildfire days.
The Poisson model is specified for each stratum using ZIP code–level death counts by calendar year and follow-up year as the outcome, with annual average all-source PM2.5 as a time-varying exposure and the log of person-years included as an offset term. We incorporate natural spline terms to model non-linear associations with PM2.5. Models are adjusted for 10 ZIP code or county-level socioeconomic and demographic covariates, four ZIP code–level meteorological variables, as well as fixed effects for geographic region (Northeast, South, Midwest, West) and calendar year to account for geographic and temporal heterogeneity. To control for individual-level variability in baseline mortality risk, each model is stratified by four individual-level characteristics: (i) a 5-year category of age at entry (65 to 69, 70 to 74, 75 to 79, 80 to 84, 85 to 89, 90 to 94, 95 to 99, and above 100 years of age), (ii) race/ethnicity (Hispanic, Native American and Alaska Native, Non-Hispanic Asian, Non-Hispanic Black, Non-Hispanic White, and other [e.g., multi-racial and missing]), (iii) sex (male or female), and (iv) an indicator variable for Medicaid eligibility (a surrogate for individual-level SES). For more details, see [wu2020evaluating].
@@
$\log\mathbb{E}[\text{death counts}]$ $\sim\text{ns(all-source PM}_{2.5}\text{, df)}$
$+\text{area-level risk factors}+\text{meteorological variables}$
$+\text{year}+\text{region}+\text{offset(log[person-years])}$
$+\text{strata(age, race, sex, Medicaid status, follow-up year)}.$
@@ Acknowledgments
Table S1: Characteristics for the Medicare study cohort, 2006–2016. Mortality and individual-level characteristics were obtained from the Centers for Medicare and Medicaid Services (CMS), and ZIP code socioeconomic status (SES) was obtained from the 2000 and 2010 Census and the 2005–2012 American Community Surveys (ACS) and county-level behavioral risk factor variables were obtained from the Centers for Disease Control and Prevention. All-source PM2.5 and wildfire PM2.5 concentrations were obtained respectively from [di2017air] and [childs2022daily]. Meteorological variables were obtained from Gridmet via Google Earth Engine.
@@
>>15% $\leq$15% Northeast Midwest South West